# ZIM: Zero-Shot Image Matting for Anything

### v1

@@ Abstract
The recent segmentation foundation model, Segment Anything Model (SAM), exhibits strong zero-shot segmentation capabilities, but it falls short in generating fine-grained precise masks. To address this limitation, we propose a novel zero-shot image matting model, called ZIM, with two key contributions: First, we develop a label converter that transforms segmentation labels into detailed matte labels, constructing the new SA1B-Matte dataset without costly manual annotations. Training SAM with this dataset enables it to generate precise matte masks while maintaining its zero-shot capability. Second, we design the zero-shot matting model equipped with a hierarchical pixel decoder to enhance mask representation, along with a prompt-aware masked attention mechanism to improve performance by enabling the model to focus on regions specified by visual prompts. We evaluate ZIM using the newly introduced MicroMat-3K test set, which contains high-quality micro-level matte labels. Experimental results show that ZIM outperforms existing methods in fine-grained mask generation and zero-shot generalization. Furthermore, we demonstrate the versatility of ZIM in various downstream tasks requiring precise masks, such as image inpainting and 3D NeRF. Our contributions provide a robust foundation for advancing zero-shot matting and its downstream applications across a wide range of computer vision tasks. The code is available at https://github.com/naver-ai/ZIM.
Figure 1: Qualitative comparison of ours with five existing zero-shot models (SAM [27], HQ-SAM [23], Matte-Any [59], Matting-Any [33], and SMat [60]). It showcases (a) box prompting results, (b) point prompting results, and (c) automatic mask generation results.
@@ 1 Introduction
In this paper, we introduce a pioneering Zero-shot Image Matting model, dubbed ZIM, that retains strong zero-shot capabilities while generating high-quality micro-level matting masks. A key challenge in this domain is the need for a matting dataset with extensive micro-level matte labels, which are costly and labor-intensive to annotate. To address this challenge, we propose a novel label conversion method that transforms any segmentation label into a detailed (pseudo) matte label. For more reliable label transformation, we design two effective strategies, $i.e.,$ Spatial Generalization Augmentation and Selective Transformation Learning, to reduce noise and yield high-fidelity matte labels (Section 3.1). Subsequently, we construct a new dataset, called SA1B-Matte, which contains an extensive set of micro-level matte labels generated by transforming segmentation labels from the SA1B dataset via the proposed converter (see Figure 2). By training SAM on the SA1B-Matte dataset, we introduce an effective foundational matting model with micro-level granularity while preserving the zero-shot ability of SAM (see Figure 1).
To further ensure effective interactive image matting, we enhance the major bottleneck in the network architecture of SAM that impedes capturing robust and detailed feature maps. Specifically, SAM employs a pixel decoder with simple two transposed convolutional layers to generate mask feature maps with a stride of 4, which is susceptible to checkerboard artifacts and often falls short in capturing fine details. To mitigate this, we implement a more elaborated pixel decoder with a hierarchical feature pyramid design, inspired by [58], enabling more robust and richer mask feature representations (Section 3.2). Furthermore, inspired by the Mask2Former [8] framework, we introduce a prompt-aware masked attention mechanism that leads to the improvement of interactive matting performance by allowing the model to focus on regions specified by visual prompts.
To validate our zero-shot matting model, we present a new test set, called MicroMat-3K, consisting of 3,000 high-quality micro-level matte labels. Our experiments on this dataset demonstrate that while SAM exhibits strong zero-shot capabilities, it struggles to deliver precise mask outputs. In contrast, existing matting models show limited zero-shot performance. ZIM, however, not only maintains robust zero-shot functionality but also provides superior precision in mask generation. Additionally, we highlight the foundational applicability of ZIM in several downstream tasks requiring precise masks, such as image inpainting [64] and 3D NeRF [6]. We hope this work provides valuable insights to the research community, encouraging further development and utilization of zero-shot matting models.
@@ 2 Related Work
Image Segmentation. Image segmentation is a fundamental task in computer vision, enabling the division of an image into distinct regions. Recent advancements in segmentation models [8, 22, 28] have significantly improved the accuracy of segmentation tasks, including semantic, instance, and panoptic segmentation. The emergence of Segment Anything Model (SAM) [27] introduced a new paradigm in segmentation by leveraging visual prompts ($e.g.,$ points or boxes). SAM is designed as a foundational segmentation model capable of handling diverse tasks due to its robust zero-shot capabilities, showing remarkable versatility across a wide range of tasks and domains. However, despite its strengths, SAM struggles to produce high-precision masks. In this paper, we address this limitation by developing a novel zero-shot model that enhances mask precision while maintaining SAM’s generalization capabilities.
@@ 3.1 Constructing the Zero-Shot Matting Dataset
Motivation. For effective zero-shot matting, a dataset with micro-level matte labels is essential. However, manually annotating matte labels at the micro-level requires extensive human labor and cost. To this end, we present an innovative Label Converter that transforms any segment label into a matte label, motivated by mask-guided matting works [41, 63]. We first collect public image matting datasets [30, 31, 29, 63, 53, 32] to train the converter. Since these datasets provide only matte labels, we derive coarse segmentation labels from matte labels by applying image processing techniques such as thresholding, resolution down-scaling, Gaussian blurring, dilation, erosion, and convex hull transformations. The converter takes an image and segmentation label as input source and is trained to produce a corresponding matte label, as illustrated in Figure 3(a).
Challenges. This approach poses two key challenges: (1) Generalization to unseen patterns. Public matting datasets predominantly contain macro-level labels ($e.g.,$ entire portraits or large object masks), as shown in Figure 2. Consequently, the converter trained on these datasets often struggles to generalize to unseen micro-level objects, such as individual body parts or detailed components. This limitation leads to the generation of noisy matte labels when applied to micro-level segmentation (see the 4th column in Figure 5(a)). (2) Unnecessary fine-grained representation. Some objects, such as cars or boxes, commonly do not require fine-grained representation. However, since the converter is trained to always transform segmentation labels into fine-grained matte labels, it often generates unnecessary noise into the output matte, particularly for objects that do not benefit from fine-grained representation (see the 4th column in Figure 5(b)). To address these challenges, we propose two simple yet effective strategies: Spatial Generalization Augmentation and Selective Transformation Learning.
@@
where $M$ and $M^{\prime}$ represent the ground-truth and predicted matte label, respectively, and $\lambda$ is a loss weighting factor hyperparameter. In addition, $\nabla_{x}$ and $\nabla_{y}$ represent the gradients along the horizontal and vertical axes, respectively. Moreover, we set a probability parameter $p$ to control the random application of Spatial Generalization Augmentation during training.
@@ 3.2 ZIM: Zero-Shot Image Matting Model
Overview of ZIM. Our proposed model, ZIM, builds upon the network architecture of SAM [27]. As illustrated in Figure 4, ZIM consists of four components: (1) Image Encoder: extracts image features from the input image, producing an image embedding with a stride of 16. (2) Prompt Encoder: encodes point or box inputs into prompt embeddings. The prompt embeddings are concatenated with learnable token embeddings, serving a role similar to the [cls] token in ViT [15]. (3) Transformer Decoder: takes the image and token embeddings to generate output token embeddings. It performs four operations: self-attention on the tokens, token-to-image cross-attention, an MLP layer, and image-to-token cross-attention that updates the image embedding. (4) Pixel Decoder: upsamples the output image embedding with a stride of 2. Lastly, the model produces matte masks by computing a dot product between the upsampled image embedding and output token embeddings.
@@
Hierarchical Pixel Decoder. To address these limitations, we introduce a hierarchical pixel decoder with a multi-level feature pyramid design, motivated by [58], as illustrated in Figure 4. The pixel decoder takes an input image and generates multi-resolution feature maps at strides 2, 4, and 8 using a series of simple convolutional layers. The image embedding is sequentially upsampled and concatenated with the corresponding feature maps at each resolution. The decoder is designed to be highly lightweight, namely adding only 10 ms of computational overhead compared to the original pixel decoder of SAM on a V100 GPU.
Our hierarchical design serves two key purposes. First, by integrating multi-level skip connections between earlier and deeper network layers, the hierarchical decoder preserves high-level semantic information while progressively refining spatial details. This combination of semantic preservation and spatial refinement results in more accurate matte outputs and reduces the potential risk of checkerboard artifacts, making the architecture more robust to challenging input prompts. Second, it enables the decoder to solidly generate high-resolution feature maps with a stride of 2, rather than the coarser stride of 4 in SAM. This finer resolution is crucial for achieving detailed matte outputs in image matting, where capturing the intricate structures of objects requires high spatial precision.
@@
Training. We train our ZIM model using the SA1B-Matte dataset. For each ground-truth matte label, we extract the corresponding box prompt from the given min-max coordinates, where the size of the box prompt is randomly perturbed up to 10% of the original size during training. Additionally, we randomly sample positive and negative point prompts following [51]. The model is optimized using the same matte loss functions defined in Eq. (1).
@@
Table 1: Quantitative comparison of our ZIM model and six existing methods on the MicroMat-3K test set. By dividing the dataset into fine-grained and coarse-grained categories, all models are evaluated on each category. Results are reported for each type of prompt (point and box) across five evaluation metrics.
@@ 4 MicroMat-3K: Zero-Shot Matting Test Set
We introduce a new test set, named MicroMat-3K, to evaluate zero-shot interactive matting models. It consists of 3,000 high-resolution images paired with micro-level matte labels, providing a comprehensive benchmark for testing various matting models under different levels of detail (see Figure 12). It includes two types of matte labels: (1) Fine-grained labels ($e.g.,$ hair, tree branches) to primarily evaluate zero-shot matting performance, where capturing intricate details is critical. (2) coarse-grained labels ($e.g.,$ cars, desks) to allow comparison with zero-shot segmentation models, which is still essential in zero-shot matting tasks. Moreover, It provides pre-defined point prompt sets for positive and negative points and box prompt sets for evaluating interactive scenarios. More detailed information about the MicroMat-3K is described in the supplementary material.
@@ 5 Experiments
Training Dataset for ZIM. To train the ZIM model, we convert the segmentation labels from the SA1B dataset [27] into matte labels using the label converter, constructing a new dataset called SA1B-Matte dataset.
Test Dataset. Models are evaluated on the MicroMat-3K, which is described in Section 4. We report results separately for fine-grained and coarse-grained object masks.
@@
Implementation Details for Label Converter. The label converter model is based on MGMatting [63] with Hiera-base-plus [49] backbone network. For training the converter, we set the input size to 1024$\times$1024, a batch size of 16, and a learning rate of 0.001 with cosine decay scheduling using the AdamW optimizer [36]. The training process runs for 500K iterations with the probability parameter $p$ for Selective Transformation Learning of 0.5 and the loss weight $\lambda$ of 10.
Implementation Details for ZIM. For the ZIM model, we use the same image encoder ($i.e.,$ ViT-B [15]) and prompt encoder as SAM. Leveraging the pre-trained weights from SAM, we fine-tune the ZIM model on 1% of the SA1B-Matte dataset, which amounts to approximately 2.2M matte labels. We set the input size to 1024$\times$1024, batch size to 16, a learning rate to 0.00001 with cosine decay scheduling using the AdamW optimizer [36], and training iterations to 500K. The loss weight $\lambda$ is set to 10 and the $\sigma$ for the point-based attention mask is set to 21 by default.
[…]
Table 2: Quantitative comparison of ZIM and existing methods on two public matting datasets, AM-2K [31] and AIM-500 [30].
@@ 5.2 Experimental Results
We evaluate ZIM against six related methods: SAM [27], SAM2 [45], HQ-SAM [23], Matte-Any [59], Matting-Any [33], and SMat [60]. All methods use the ViT-B [15] backbone network, except SAM2, which uses Hiera-base-plus [49]. Table 1 presents the evaluation scores across five metrics for both point and box prompts on fine-grained and coarse-grained masks. For coarse-grained results, SAM achieves reasonable zero-shot performance, while other methods ($e.g.,$ Matting-Any and SMat) struggle to generalize to unseen objects. This is likely due to their fine-tuning on macro-level labeled datasets, which degrades their zero-shot capabilities. The qualitative results in Figure 1 show that these methods often produce macro-level outputs, even when provided with micro-level prompts. Moreover, SAM tends to suffer from checkerboard artifacts when challenging prompts are introduced. In contrast, ZIM generates more robust and higher-quality masks, due to our hierarchical feature pyramid decoder, as shown in Figure 1. The fine-grained results in Table 1 highlight the superiority of our method in producing high-quality matting outputs while maintaining strong zero-shot capabilities.
In addition, Table 2 provides quantitative results on two public matting datasets ($i.e.,$ AM-2K [31] and AIM [30]), which contain macro-level matte labels only. Here, existing matting methods ($e.g.,$ Matte-Any, Matting-Any, and SMat) outperform SAM, as they are specifically fine-tuned on macro-level labeled training datasets that are closely aligned with the domain of the test set. Despite being trained on the micro-level SA1B-Matte dataset, ZIM also achieves highly competitive performance on these matting datasets, showing superior generalization capability.
Lastly, Table 7 provides an in-depth analysis, including throughput, backbone network effects, and detailed scores based on object size. This analysis demonstrates that ZIM’s components (hierarchical mask decoder and prompt-aware masked attention) are lightweight, adding only 10 ms more throughput compared to SAM, while making ZIM more robust to checkerboard artifacts, particularly in larger objects.
Figure 5: Qualitative analysis of key components of the label converter: qualitative samples (a) without Spatial Generalization Augmentation (SGA) and (b) without Selective Transformation Learning (STL).

### v2

@@ Abstract
The recent segmentation foundation model, Segment Anything Model (SAM), exhibits strong zero-shot segmentation capabilities, but it falls short in generating fine-grained precise masks. To address this limitation, we propose a novel zero-shot image matting model, called ZIM, with two key contributions: First, we develop a label converter that transforms segmentation labels into detailed matte labels, constructing the new SA1B-Matte dataset without costly manual annotations. Training SAM with this dataset enables it to generate precise matte masks while maintaining its zero-shot capability. Second, we design the zero-shot matting model equipped with a hierarchical pixel decoder to enhance mask representation, along with a prompt-aware masked attention mechanism to improve performance by enabling the model to focus on regions specified by visual prompts. We evaluate ZIM using the newly introduced MicroMat-3K test set, which contains high-quality micro-level matte labels. Experimental results show that ZIM outperforms existing methods in fine-grained mask generation and zero-shot generalization. Furthermore, we demonstrate the versatility of ZIM in various downstream tasks requiring precise masks, such as image inpainting and 3D segmentation. Our contributions provide a robust foundation for advancing zero-shot matting and its downstream applications across a wide range of computer vision tasks. The code is available at https://naver-ai.github.io/ZIM.
Figure 1: Qualitative comparison of ours with five existing zero-shot models (SAM [34], HQ-SAM [29], Matte-Any [79], and Matting-Any [41]). It showcases (a) box prompting results, (b) point prompting results, and (c) automatic mask generation results.
@@ 1 Introduction
In this paper, we introduce a pioneering Zero-shot Image Matting model, dubbed ZIM, that retains strong zero-shot capabilities while generating high-quality micro-level matting masks. A key challenge in this domain is the need for a matting dataset with extensive micro-level matte labels, which are costly and labor-intensive to annotate. To address this challenge, we propose a novel label conversion method that transforms any segmentation label into a detailed matte label. For more reliable label transformation, we design two effective strategies to reduce noise and yield high-fidelity matte labels (Section 3.1). Subsequently, we construct a new dataset, called SA1B-Matte, which contains an extensive set of micro-level matte labels generated by transforming segmentation labels from the SA1B dataset via the proposed converter (see Figure 2). By training SAM on the SA1B-Matte dataset, we introduce an effective foundational matting model with micro-level granularity while preserving the zero-shot ability of SAM (see Figure 1).
To further ensure effective image matting, we enhance the major bottleneck in the network architecture of SAM that impedes capturing robust and detailed feature maps. Specifically, SAM employs a simple pixel decoder to generate mask feature maps with a stride of 4, which is susceptible to checkerboard artifacts and often falls short in capturing fine details. To mitigate this, we design a more elaborated pixel decoder, enabling more robust and richer mask representations (Section 3.2). Furthermore, we introduce a prompt-aware masked attention mechanism that leads to the improvement of interactive matting performance.
To validate our zero-shot matting model, we present a new test set, called MicroMat-3K, consisting of 3,000 high-quality micro-level matte labels. Our experiments on this dataset demonstrate that while SAM exhibits strong zero-shot capabilities, it struggles to deliver precise mask outputs. In contrast, existing matting models show limited zero-shot performance. ZIM, however, not only maintains robust zero-shot functionality but also provides superior precision in mask generation. Additionally, we highlight the foundational applicability of ZIM in several downstream tasks requiring precise masks, such as image inpainting [84] and 3D segmentation [8]. We hope this work provides valuable insights to the research community, encouraging further development of zero-shot matting models.
@@ 2 Related Work
Image Segmentation. Image segmentation is a fundamental task in computer vision, enabling the division of an image into distinct regions. Recent advancements in segmentation models [12, 28, 35] have significantly improved the accuracy of segmentation tasks, including semantic and instance segmentation. The emergence of Segment Anything Model (SAM) [34] introduced a new paradigm in segmentation by leveraging visual prompts ($e.g.,$ points or boxes). SAM is designed as a foundational segmentation model capable of handling diverse tasks due to its robust zero-shot capabilities, showing remarkable versatility across a wide range of tasks and domains. However, despite its strengths, SAM struggles to produce high-precision masks. In this paper, we address this limitation by developing a novel zero-shot model that enhances mask precision while maintaining SAM’s generalization capabilities.
@@ 3.1 Constructing the Zero-Shot Matting Dataset
Motivation. For effective zero-shot matting, a dataset with micro-level matte labels is essential. However, manually annotating matte labels at the micro-level requires extensive human labor and cost. To this end, we present an innovative Label Converter that transforms any segment label into a matte label, motivated by mask-guided matting works [55, 83]. We first collect public image matting datasets [37, 38, 36, 83, 69, 40] to train the converter. We derive coarse segmentation labels from matte labels by applying image processing techniques such as thresholding, resolution down-scaling, Gaussian blurring, dilation, erosion, and convex hull transformations. The converter takes an image and segmentation label as input source and is trained to produce a corresponding matte label, as illustrated in Figure 3(a).
Challenges. (1) Generalization to unseen patterns: Public matting datasets predominantly contain macro-level labels ($e.g.,$ entire portraits), as shown in Figure 2. Consequently, the converter trained on these datasets often struggles to generalize to unseen micro-level objects ($e.g.,$ individual body parts). This limitation leads to the generation of noisy matte labels when applied to micro-level segmentation (see the 4th column in Figure 6(a)). (2) Unnecessary fine-grained representation: Some objects, such as cars or boxes, commonly do not require fine-grained representation. However, since the converter is trained to always transform segmentation labels into fine-grained matte labels, it often generates unnecessary noise into the output matte, particularly for objects that do not benefit from fine-grained representation (see the 4th column in Figure 6(b)).
@@
where $M$ and $M^{\prime}$ represent the ground-truth and predicted matte label, respectively, and $\lambda$ is a loss weighting factor. In addition, $\nabla_{x}$ and $\nabla_{y}$ represent the gradients along the horizontal and vertical axes, respectively. Moreover, we set a probability parameter $p$ to control the random application of Spatial Generalization Augmentation during training.
@@ 3.2 ZIM: Zero-Shot Image Matting Model
Overview of ZIM. Our proposed model, ZIM, builds upon SAM [34] and consists of four components, as illustrated in Figure 4: (1) Image Encoder: extracts image features from the input image, producing an image embedding with a stride of 16. (2) Prompt Encoder: encodes point or box inputs into prompt embeddings concatenated with learnable token embeddings, serving a role similar to the [cls] token in ViT [20]. (3) Transformer Decoder: takes the image and token embeddings to generate output token embeddings. It performs four operations: self-attention on the tokens, token-to-image cross-attention, an MLP layer, and image-to-token cross-attention that updates the image embedding. (4) Pixel Decoder: upsamples the output image embedding with a stride of 2. Lastly, the model produces matte masks by computing a dot product between the upsampled image embedding and output token embeddings.
@@
Hierarchical Pixel Decoder. To address these shortcomings, we introduce a hierarchical pixel decoder with a multi-level feature pyramid design, motivated by [78], as illustrated in Figure 4. The pixel decoder takes an input image and generates multi-resolution feature maps at strides 2, 4, and 8 using a series of simple convolutional layers. The image embedding is sequentially upsampled and concatenated with the corresponding feature maps at each resolution. The decoder is designed to be highly lightweight, namely adding only 10 ms of computational overhead compared to the original pixel decoder of SAM on a V100 GPU. Our hierarchical design serves two key purposes: First, it preserves high-level semantics while refining spatial details, reducing checkerboard artifacts and enhancing robustness to challenging prompts. Second, it generates high-resolution feature maps with a stride of 2, essential for capturing fine-grained structures in matting.
@@
Training. We train ZIM using the SA1B-Matte dataset. From the ground-truth matte label, we extract a box prompt from the given min-max coordinates and randomly sample positive and negative point prompts following [66]. The model is optimized using the same matte loss functions defined in Eq. (1).
@@
Table 1: Quantitative comparison of ZIM and six existing methods on the MicroMat-3K test set, evaluated separately on fine-grained and coarse-grained categories using point and box prompts across five evaluation metrics.
@@ 4 MicroMat-3K: Zero-Shot Matting Test Set
We introduce a new test set, named MicroMat-3K, to evaluate zero-shot interactive matting models. It consists of 3,000 high-resolution images paired with micro-level matte labels. It includes two types of matte labels: (1) Fine-grained labels ($e.g.,$ hair, tree branches) to primarily evaluate zero-shot matting performance, where capturing intricate details is critical. (2) coarse-grained labels ($e.g.,$ cars, desks) to allow comparison with zero-shot segmentation models, which is still essential in zero-shot matting tasks. Moreover, It provides pre-defined point prompt sets for positive and negative points and box prompt sets for evaluating interactive scenarios. More detailed information about the MicroMat-3K is described in the supplementary material.
@@ 5 Experiments
Figure 5: Downstream Transferability Evaluation by replacing SAM with ZIM as the segmentation foundation model in various downstream tasks: Matte-Any [79] and Matting-Any [41] are evaluated on MicroMat3K-Fine, AIM-500 [37], AM-2K [38], and P3M-500-NP [36] using the MSE metric. HQ-SAM [29] is assessed on DIS [59], COIFT [44], HRSOD [85], and ThinObject [44] using mIoU. Inpainting Anything [84] is tested on COCO [45] using CLIP distance. Following Medical Image Segmentation evaluation protocol [50] with five different prompting modes (mIoU). 3D segmentation follows the SA3D [8] framework and is evaluated on NVOS [62].
Implementation Details for Label Converter. The label converter model is based on MGMatting [83] with Hiera-base-plus [64] backbone network. For training the converter, we set the input size to $\times1024$, a batch size of 16, and a learning rate of 0.001 with cosine decay scheduling using the AdamW optimizer [48]. The training process runs for 500K iterations with the probability parameter $p$ of 0.5 and the loss weight $\lambda$ of 10.
Implementation Details for ZIM. For the ZIM model, we use the same image encoder ($i.e.,$ ViT-B [20]) and prompt encoder as SAM. Leveraging the pre-trained weights from SAM, we fine-tune the ZIM model on 1% of the SA1B-Matte dataset, which amounts to approximately 2.2M matte labels. We set the input size to $\times1024$, batch size to 16, a learning rate to 0.00001 with cosine decay scheduling using the AdamW optimizer [48], and training iterations to 500K. The loss weight $\lambda$ is set to 10 and the $\sigma$ for the point-based attention mask is set to 21 by default.
@@ 5.2 Experimental Results
We evaluate ZIM against four related methods on MicroMat3K: SAM [34], HQ-SAM [29], Matte-Any [79], and Matting-Any [41]. All methods use the ViT-B [20] backbone network. Table 1 presents the evaluation scores across five metrics for both point and box prompts on fine-grained and coarse-grained masks. For coarse-grained results, SAM achieves reasonable zero-shot performance, while other methods ($e.g.,$ Matting-Any and HQ-SAM) struggle to generalize to unseen objects. This is likely due to their fine-tuning on macro-level labeled datasets, which degrades their zero-shot capabilities. The qualitative results in Figure 1 show that these methods often produce macro-level outputs, even provided with micro-level prompts. Moreover, SAM tends to suffer from checkerboard artifacts when challenging prompts are introduced. In contrast, ZIM generates a more robust quality of masks, due to our hierarchical feature pyramid decoder. The fine-grained results in Table 1 highlight ZIM’s superiority in producing high-quality matting outputs while maintaining strong zero-shot capabilities.
5.3 Downstream Transferability
We assess the transferability of ZIM compared to SAM across various downstream tasks. Specifically, we first integrate ZIM into existing matting methods, Matte-Any [79] and Matting-Any [41], and observe significantly improved box prompting MSE results across diverse matting benchmarks, including MicroMat3K-Fine, AIM-500 [37], AM-2K [38], and P3M-500-NP [36]. Likewise, replacing SAM with ZIM in HQ-SAM [29] notably enhances performance in fine-grained segmentation datasets, including DIS [59], COIFT [44], HRSOD [85], and ThinObject [44].
Moreover, we extend our analysis to broader vision tasks, including image inpainting, medical image segmentation, and 3D segmentation. Specifically, we evaluate image inpainting using the Inpainting Anything framework [84] on the COCO dataset [45] via the CLIP Distance metric [21, 81]. For medical image segmentation, we utilize the zero-shot evaluation protocol from [50] on five diverse medical imaging datasets [2, 67, 53, 25], measured by mean IoU (mIoU). Additionally, we assess 3D segmentation performance within the SA3D framework [8] on the NVOS dataset [62]. These tasks inherently require highly precise segmentation masks for optimal outcomes. While replacing SAM with existing matting models ($e.g.,$ Matte-Any and Matting-Any) results in significant performance drops due to limited generalization capabilities, employing ZIM consistently boosts zero-shot performance across these diverse scenarios, highlighting its robust generalization and precise mask representation capabilities. Comprehensive qualitative and quantitative results for these tasks are provided in the supplementary materials.
Figure 6: Qualitative analysis of key components of the label converter: (a) without Spatial Generalization Augmentation (SGA) and (b) without Selective Transformation Learning (STL).